OpenStudy (anonymous):

Find the common ratio of the geometric sequence: 1.2, 0.84, 0.588, 0.4116

OpenStudy (anonymous):

Common ratio = .84 /1.2 = 7/10

OpenStudy (anonymous):

or 0.4116/0.588= 7/10

OpenStudy (anonymous):

Common Ratio is always a constant. common ratio will give the same result in any case second term / first term,third term /second term ,fourth term/third term .....

OpenStudy (anonymous):

0.4116=1.2*r^(4-1) 0.4116/1.2 =r^3\[r=\sqrt[3]{0.343}=0.7\]

OpenStudy (anonymous):

No need to apply tn = a r^(1-n) r you can get directly as i said

OpenStudy (anonymous):

many way to solve

OpenStudy (anonymous):

I know this is offtopic, but is this advanced geometry or calculus 2?